For a crystal, the refractive index no for the o-ray is 1.5442 for light of wavelength 

6×10-5 m. The least thickness of the crystal used as a quarter wave plate is found to 

be 1.65×10-5 m. Determine the refractive index ne for the e-ray in this crystal.



1
Expert's answer
2021-06-03T18:26:28-0400

"(n_e-n_o)d=\frac{\lambda}{4},"

"n_e=n_o+\frac{\lambda}{4d},"

"n_e=1.5442+\frac{6}{4\cdot 1.65}=2.4533."